# Prompt Selection Matters: Enhancing Text Annotations for Social Sciences with Large Language Models

### v1

@@ 5 Conclusion
This paper meant to illustrate the importance of prompt selection and the effectiveness of automatic prompt optimization within the context of text annotation tasks frequently used in the social sciences. In particular, our results suggest that the prompt optimization procedure described in Subsection 3.3 can be applied off-the-shelf to various tasks to achieve accuracies that equal or exceed those obtained using hand-crafted prompts. This prompt optimization method is made particularly easy by our user-friendly implementation888In particular, it allows for the replication of all of our experiments without the need for any coding., which is freely available at https://prompt-ultra.github.io/.
@@
Among other potential future research directions, it would be particularly interesting to test whether LLMs tasked with labelling a corpus can give robust justifications for their choices – e.g. explain that they classified a given tweet as left-wing due to the presence of such and such opinions that are typically associated with the Left. Though more sophisticated solutions are conceivable (e.g. using the LLMs’ attention mechanisms, see [VSP+17]), directly asking the chatbot "Can you justify your decision?" would already make for an interesting experiment. What would also be extremely valuable is to be able to ask the LLMs to associate a (reliable) confidence score to its labels, so that human annotators can review the small percentage of labels the model was unsure of, and increase accuracy.
Appendix A Brief overview of our browser-based service
We summarily describe our automatic text labelling service https://prompt-ultra.github.io/.

### v2

@@ 6 Conclusion
This paper meant to illustrate the importance of prompt selection and the effectiveness of automatic prompt optimization within the context of text annotation tasks frequently used in the social sciences. In particular, our results suggest that the prompt optimization procedure described in Subsection 4.3 can be applied off-the-shelf to various tasks to achieve accuracies that equal or exceed those obtained using hand-crafted prompts. This prompt optimization method is made particularly easy by our user-friendly implementation777In particular, it allows for the replication of all of our experiments without the need for any coding., which is freely available at https://prompt-ultra.github.io/.
@@
Among other potential future research directions, it would be particularly interesting to test whether LLMs tasked with labelling a corpus can give robust justifications for their choices – e.g. explain that they classified a given tweet as left-wing due to the presence of such and such opinions that are typically associated with the Left. Though more sophisticated solutions are conceivable (e.g. using the LLMs’ attention mechanisms, see [VSP+17]), directly asking the chatbot "Can you justify your decision?" would already make for an interesting experiment. Getting LLMs to associate a (reliable) confidence score to its labels, so that human annotators can review the small percentage of labels the model was unsure of and increase accuracy, would also be of great interest.
7 Data availability statement
All data used in our analyses are linked in the "3.1 Datasets and tasks" section of the manuscript.
8 Conflict of interest
Nothing declared.
Appendix A Results of the main experiments as numerical tables
Micro averaged F1-scores | ||||
TE-hate | TE-emotion | TE-sent | TE-off | |
Simple | 57.0 [55.2, 58.8] | 60.5 [59.5, 61.4] | 71.4 [68.4, 74.4] | |
Explanations | ||||
Examples | 76.9 [74.7, 79.1] | |||
Roleplay | ||||
CoT | ||||
APO | ||||
TML-sent | AS-pol | LibCon | ||
Simple | ||||
Explanations | ||||
Examples | ||||
Roleplay | 65.5 [64.4, 66.7] | 47.1 [46.1, 48.1] | 70.7 [69.8, 71.5] | |
CoT | ||||
APO |
Train set vs test set | ||||
TE-hate, train | TE-hate, test | TE-emotion, train | TE-emotion, test | |
Simple | ||||
Explanations | ||||
Examples | ||||
Roleplay | ||||
CoT |
@@ Appendix B Optimized prompts
Appendix C Brief overview of our browser-based service
We summarily describe our automatic text labelling service https://prompt-ultra.github.io/. Further explanations are given on the website itself.